OpenStudy (anonymous):

help with inverse? find inverse of A(t)= 75-14ln(1+t)?

OpenStudy (anonymous):

14ln(1+t) = 75 - A ln(1+t) = (75-A)/ 14 1+ t = e ^(75-A)/ 14 t = e ^(75-A)/ 14 - 1 inverse A(t)-1= [ e ^(75-t)/ 14) ]- 1

OpenStudy (anonymous):

along with that I am confused on how to find t if A(t)=50?

OpenStudy (anonymous):

ok use t = e ^(75-A)/ 14) - 1 = e^((75-50)/ 14) - 1 = e^(25/14) - 1 = 4.964 to 3 dec. places

OpenStudy (anonymous):

ohh okay thats why i didnt get that part. it makes sense now

OpenStudy (anonymous):

fine